OpenStudy (anonymous):

can someone pls explain how we get: (delta) angular displacement (theta) = (delta) V/ v

OpenStudy (irishboy123):

\[\Delta \theta = \frac{\Delta V}{v}\] ??? is there a context?

OpenStudy (anonymous):

YEP THIS IS THE DERIVATION TO FIND OUT a = v^2 / r

OpenStudy (irishboy123):

what is big V and small v in your question? as it happens, i posted a derivation on another thread today http://openstudy.com/study#/updates/55fbfc07e4b0027d1e7739c7